# Value Profiles for Encoding Human Variation

### v1

@@ 6 Value Profile Clustering for Grouping Raters
For DIC and OQA, clustering is highly effective - using just two clusters preserves the majority of the usable rater information (60%/51% respectively), and using eight clusters roughly matches the performance of giving each rater their own profile. This suggests that perhaps most raters fall into one of very few “modes” of agreement for these datasets, and that clustering based on value profiles is highly effective at finding these groupings. For the other two large datasets, HK and VP, clustering preserves a significant amount of information ($\geq 20\%$) but is not as predictive. This implies either a failure to find the best clusters or that the underlying variation is inherently more difficult to categorize. Interestingly, this dataset divide coheres with our intuitions: e.g., for OpinionQA, ideology is highly explanatory and mostly centered around a few clusters, whereas ValuePrism includes a diverse set of 4k unique values which resist categorization. While it is epistemically difficult to totally disentangle a failure of our method to find correct groupings vs. a true difference in dimensionality of disagreement, we do find these results suggestive of profile clustering being able to tell us something interesting about the true underlying reasons for rater variation.
Profile clusters are more predictive than the best demographic groupings. Next, we compare with the best performing demographic clusters, grouping people who gave the same demographic response to a categorical demographic question (e.g., people in the same country for DIC or same political ideology for OQA). We compare specifically across the three large datasets w/ demographics: for DIC, the two profile clusters is more predictive than grouping based on country; for HK/OQA, the four profile clusters outperform grouping by religiosity/ideology respectively. In other words, clustering by value profiles is able to outperform the most performant unidimensional demographic clusters when using the same or fewer number of groupings.
Cluster descriptions qualitatively describe modes of disagreement.. The profile clustering algorithm returns not only clustering assignments, but also a single corresponding value profile for each cluster (see App. H). For DICES, even two clusters were quite predictive. The corresponding value clusters relate to overall sensitivity to toxicity (e.g., profile 1: ”High tolerance for offensive language”; ”Narrow definition of toxicity” vs. profile 2: ”Strong reaction to overt negativity, ”Sensitivity to potential harm”). Meanwhile, when going to four clusters, more nuance enters in (e.g., ”Context and intent matter”). In other words, it seems that 1) overall sensitivity to toxicity is an important dimension in explaining variation, and 2) there are clusters of people who hold more nuanced views. For OpinionQA, descriptors that have to do with politics are often used (e.g., ”Economically Conservative, but Populist on Trade”). For HK, which required more clusters to effectively predict, there are many specific phrases about what kinds of things the rater may or may not find offensive (e.g., ”Profanity tolerance”; ”Discomfort with stereotyping”; ”Acceptance of strong opinions”; etc.). Meanwhile, for PR the clusters center around potentially conflicting chatbot preferences (e.g., ”Appreciates simplicity” vs. ”Appreciates nuanced and comprehensive answers”). Because value profiles are interpretable (see §7) and can recover demographic groupings where predictive, we have reason to believe that these qualitative differences map to important dimensions of disagreement for a dataset.
7 Extrinsic Evaluation
Value Profiles are interpretable. We first explore interpretability – in other words, do the value descriptions change the decoder outputs in a common-sense manner? Because the encoder is prompted and only the decoder is trained (cf. Section 2.2), we believe that this serves as a strong regularization so that the value profiles correlate with held-out ratings only by the natural language values described. To ensure that this is the case, we test the autoencoder for interpretability as follows: 1. For each instance and 100 value profiles, we get the estimated output distribution for the decoder. 2. We select the value profiles that have the largest Jensen-Shannon divergence. 3. We create the following binary classification task: given an instance and two value profiles, which profile corresponds to which estimated output distribution? If the change in distribution is not correlated with common sense, we would expect 50% accuracy, while performance would approach 100% if an observer is always able to tell which value profile goes with which distribution.
Data Accuracy (%)
OQA 94.8 (±$\pm 2.5$)
HK 96.3 (±.$\pm.6$)
DIC 95.5 (±.$\pm.7$)
VP 91.7 (±.$\pm.9$)
HL 89.8 (±.$\pm.9$)
PR 80.0 (±.$\pm.3$)
Chance 50
Table 2: Profiles are semantically interpretable (95% CI).
As an example of the task, here is a shortened example from VP: Is it moral, immoral, or morally dependent on context to do this: “Choosing not to get a vaccine.” Profile 1: Prioritization of collective good over individual needs. Profile 2: Strong belief in individual liberty and autonomy. X probabilities: Moral: 92%, Context-dependent: 7%, Immoral: 2%. Y probabilities: Immoral: 94%, Context-dependent: 3%, Moral: 3%. Which profile goes with the X probabilities? Correct answer in footnote.  [^5] Profile two is the correct answer.
We report accuracies for a zero-shot prompted Gemini in Table 2. Accuracies range from 80-96% across all datasets, demonstrating that variation in outputs from value profiles are explainable by their plain natural language (p$<$.001 for all datasets).
Figure 10: Calibration plots for value profile decoders. (Perfect calibration $=$ dotted line). The decoders are very well-calibrated.
Decoders are well-calibrated. Decoder calibration is important for two principal reasons. Firstly, an appropriately calibrated decoder allows us to trust the model confidence w.r.t. error rate. Secondly, even raters with shared values may have varied outputs - a well-trained decoder would model this distribution appropriately. In other words, calibration is important for epistemic uncertainty of the decoder and the aleatoric uncertainty of the population who holds those values. Calibration plots for the value profile decoders can be found in Figure 10666 We observe similarly well-calibrated decoders for other rater representations. The trained decoders are quite well-calibrated, suggesting that we can generally trust the decoder’s output confidence. Proper calibration allows for exciting potential applications, such as disentangling value-related epistemic uncertanity from aleatoric uncertainty in rater variation. (See App. D for more discussion.)
Figure 11: Instance-level observed vs. estimated interannotator agreement (as the probability that two raters agree). The predicted simulated agreement correlates with the observed agreement.
Simulating an annotator population with value profiles. Given a trained decoder and a set of value profiles, one can simulate a population – or “jury” (Gordon et al., 2022) – of annotators on novel instances. While one can do many things with such a simulated population (Park et al., 2023; Gordon et al., 2022), one experiment is to predict which instances raters would have higher or lower inter-annotator agreement (IAA).
[…]
8 Related Work
Clustering and demographics While aligning to groups can increase agreement (Chen et al., 2024), it also has been shown to flatten intra-group variation (Orlikowski et al., 2023; Wang et al., 2025) lead to stereotyping (Cheng et al., 2023), or simply not be correlated with subjective NLP tasks Orlikowski et al. (2025). Prior work proposes clustering individuals by responses (Vitsakis et al., 2024; Li et al., 2024) and similarly finds that clusters cut across demographic groups. Additionally, unlike these methods, our value profile clustering method does not require that annotators label the same instances. Beyond predictivity, demographics can still be important to collect for evaluating group fairness (Aguirre et al., 2023; Kirk et al., 2024b).
Steering to individuals Prompted large language models (LLMs) to simulate human judgments such as: NLP task annotators (Bavaresco et al., 2024), political survey respondents (Argyle et al., 2023), fact-checking labels (De et al., 2024), or human attitudes and behavior (Park et al., 2024). Textual user profiles have also been proposed for personalizing chats (Zhang et al., 2018). Many works focus merely on prompting an existing LLM, while our work explicitly trains an LLM to better match varied perspectives (Gordon et al., 2022; Jiang et al., 2024). Encoding individual information from demonstrations is also analogous to behavioral user modeling for recommender systems (Balog et al., 2019; Radlinski et al., 2022; Ramos et al., 2024).
Values and alignment Similarly to natural language value profiles, Bai et al. (2022)’s ”constitutional AI” train models to follow textual principles, although they focus on a single set of principles. Findeis et al. (2024) propose to learn preference principles directly from preference data (similar to our encoder setup). Values have also been normatively proposed as an alignment target (Gabriel, 2020; Klingefjord et al., 2024), and pluralistic alignment (Sorensen et al., 2024b) seeks to align AI systems to diverse values.
9 Conclusion
10 Ethical Considerations
Profiling risks One of the potential ethical risks of the value profile through an autoencoder paradigm is in the name: “profile.” Value profiles are, inherently, guesses about the underlying values that people may hold that lead them to annotate in the way that they do. There are potential privacy concerns here - people may not wish to have their underlying values exposed (Tomasev et al., 2021). It may be better if people had agency to create their own value profiles through some elicitation process, such as qualitative interviews (Park et al., 2024), such that people can choose exactly what value-relevant information to expose.
False generalization On the one hand, value profiles are an attempt to reduce the (often false) generalization inherent when grouping e.g. by demographic groups Dev et al. (2022), improving on widely used current techniques. On the other hand, generalization risks remain. For example, there may be many be multiple possible underlying values that could support a set of rater annotations. In the absence of additional information, the value encoder may arbitrarily assign a guess to the underlying values. Also, our experiments focus on English-language value profiles, so generalization to other languages is unknown. Also, if people hold values in ways that are not commonly understood/represented by natural language, we would expect performance degradation. may lead to false generalization about a rater. Additionally, there is always a risk of misrepresentation when using simulated human ratings in place of actual ratings at all (Agnew et al., 2024).
Interpretability, understandability, and user agency However, there are also several positive attributes to value profiles. First of all, they are interpretable - and therefore, potentially more understandable to a user. While people interact with many technologies today that are trying to model their behavior and preferences, most such systems do not break down their user representation into a format that as as easy to understand as a textual description. Additionally, this makes value profiles intervenable - people could change how they choose to be represented. For example, one could imagine that after a value profile is inferred for a person, they would have the chance to change the value profile, as suggested by Balog et al. (2019); Lazar et al. (2024) for personalized recommender systems. Relatedly, value profiles serve as a step towards explainable AI (Arrieta et al., 2019; Koh et al., 2020) for human variation.
Enabling value reflection Learning values from data, while allowing users to modify the values, is loosely related to John Rawls’ concept of reflective equilibrium (Rawls, 2005; Knight, 2025): ratings are akin to judgments, and value profiles are an attempt to draw general “principles” out of the judgments in a bottom-up manner. Meanwhile, a user can then edit the value profile, applying top-down reflection on whether the values/principles are ones that the person would endorse. In this light, perhaps “value profiles” could help a person to explore their own value system, both in the values that their decisions seem to imply that they may have and considering which values they would reflexively endorse.
”Chosenness” of values Many works modeling diversity focus on socio-demographics. However, many demographics are not a result of a person’s agency, but rather a product of unchosen life factors - for example, the country in which one is born, or the economic opportunities available to them. Meanwhile, while the values that one holds can certainly be affected by unchosen factors (Nguyen, 2024), values can also be chosen for oneself. Thus, in the spirit of luck egalitarianism (Dworkin, 2002; Lippert-Rasmussen, 2023) it may be more justifiable to represent someone using values that they reflexively endorse, as opposed to boxing them in to the characteristics of a group they may not have chosen.
11 Reproducibility Statement
@@
$N_{x}\leftarrow|\{x_{j}\text{ s.t. }\exists i,(x_{j},\cdot)\in R_{i}^{\text{fit}}\}|$triangleright # unique inst.
Initialize $P\in\mathbb{R}^{N_{x}\times N_{v}\times|\mathcal{Y}|}$triangleright Fill in output probabilities conditioned on each value profile
@@
         $P[j,k]=d(x_{j},v_{k})$triangleright Prob. dist. over $\mathcal{Y}$ from $d$ conditioned on instance $x_{j}$ and profile $v_{k}$
@@
$L(r_{i},v_{k})\leftarrow\sum_{(x_{j},y_{ij})\in R_{i}^{\text{fit}}}-\log P[j,k,y_{ij}]$triangleright Total loss from assigning rater $r_{i}$ to profile $v_{k}$
// Approximately find the $N_{\textrm{cluster}}$ value profiles such that loss is minimized when assigning each rater to one of the profiles
// e.g., $\underset{\hat{C}\in\mathcal{P}(V):|\hat{C}|=N_{\textrm{cluster}}}{\arg\min}\sum_{r_{i}\in R_{i}^{\textrm{fit}}}\min_{\tilde{v}\in\hat{C}}L(r_{i},\tilde{v})$
@@
         $C[c]\leftarrow$triangleright New cluster that minimizes loss
         $\underset{\hat{v}_{c}\in V}{\arg\min}\underset{i\in[1,N_{\text{R}}]}{\sum}\underset{\tilde{v}\in(C/\{\hat{v}_{c}\}\cup\{\tilde{v}_{c}\})}{\min}L(r_{i},\tilde{v})$
@@
Output: Clusters C, assignments $\underset{\tilde{v}\in C}{\arg\min}L(r_{i},\tilde{v})$
@@
Note: The most expensive portion in the clustering algorithm is running the $N_{V}\times N_{X}$ decoder inferences - finding the clusters only requires lookups in table $P$. Thus, the biggest determiner of runtime is the number of unique instances $N_{X}$ and the number of value profiles considered $N_{V}$.
Decoder hyperparameters: model: gemma2-9b-pt (Team et al., 2024b), batch size: 4, learning rate: 1e-7, gradient clipping: 50.
fp32 unembedding layer: Gemma 2 (Team et al., 2024b) natively uses bf16. However, we found that this caused heavy quantization among high-probability logits (e.g., the valid responses). As such, we cast the embedding/unembedding parameters to fp32 before training, which allowed for higher precision distributions, important for calibration and expressivity.
Fit/eval partition details: For each rater $r_{i}$, we draw $|\mathcal{D}_{i}^{\text{fit}}|\sim\mathcal{U}(\{2,\ldots,|\mathcal{D}_{i}|-2\})$ and set $|\mathcal{D}_{i}^{\text{eval}}|=|\mathcal{D}_{i}|-|\mathcal{D}_{i}^{\text{eval}}|$ to ensure that we have variable-sized fit/eval splits with at least two instances each. Value profile encoders use all $D_{i}^{\text{fit}}$ instances and the decoders with in-context information $E_{n}$ use the first $\min(n,|D_{i}^{\text{fit}}|)$ examples from $|D_{i}^{\text{fit}}|$. This means that value profiles are fit with a variable number of ratings.

### v2

@@ 6 Value Profile Clustering for Grouping Raters
For DIC and OQA, clustering is highly effective - using just two clusters preserves the majority of the usable rater information (60%/51% respectively), and using eight clusters roughly matches the performance of giving each rater their own profile. This suggests that perhaps most raters fall into one of very few “modes" of agreement for these datasets, and that clustering based on value profiles is highly effective at finding these groupings. For the other two large datasets, HK and VP, clustering preserves a significant amount of information ($\geq 20\%$) but is not as predictive. This implies either a failure to find the best clusters or that the underlying variation is inherently more difficult to categorize. Interestingly, this dataset divide coheres with our intuitions: e.g., for OpinionQA, ideology is highly explanatory and mostly centered around a few clusters, whereas ValuePrism includes a diverse set of 4k unique values which resist categorization. While it is epistemically difficult to totally disentangle a failure of our method to find correct groupings vs. a true difference in dimensionality of disagreement, we do find these results suggestive of profile clustering being able to tell us something interesting about the true underlying reasons for rater variation.
@@
$N_{x}\leftarrow|\{x_{j}\text{ s.t. }\exists i,(x_{j},\cdot)\in R_{i}^{\text{fit}}\}|$ $\triangleright$ # unique inst.
Initialize $P\in\mathbb{R}^{N_{x}\times N_{v}\times|\mathcal{Y}|}$ $\triangleright$ Fill in output probabilities conditioned on each value profile
@@
   $P[j,k]=d(x_{j},v_{k})$ $\triangleright$ Prob. dist. over $\mathcal{Y}$ from $d$ conditioned on instance $x_{j}$ and profile $v_{k}$
@@
$L(r_{i},v_{k})\leftarrow\sum_{(x_{j},y_{ij})\in R_{i}^{\text{fit}}}-\log P[j,k,y_{ij}]$ $\triangleright$ Total loss from assigning rater $r_{i}$ to profile $v_{k}$
@@
   $C[c]\leftarrow$ $\triangleright$ New cluster that minimizes loss
[…]
@@
[…]
@@
Profile clusters are more predictive than the best demographic groupings. Next, we compare with the best performing demographic clusters, grouping people who gave the same demographic response to a categorical demographic question (e.g., people in the same country for DIC or same political ideology for OQA). We compare specifically across the three large datasets w/ demographics: for DIC, the two profile clusters is more predictive than grouping based on country; for HK/OQA, the four profile clusters outperform grouping by religiosity/ideology respectively. In other words, clustering by value profiles is able to outperform the most performant demographic clusters when using the same or fewer number of groupings.
Cluster descriptions qualitatively describe modes of disagreement.. The profile clustering algorithm returns not only clustering assignments, but also a single corresponding value profile for each cluster (see App. H). For DICES, even two clusters were quite predictive. The corresponding value clusters relate to overall sensitivity to toxicity (e.g., profile 1: "High tolerance for offensive language"; "Narrow definition of toxicity" vs. profile 2: "Strong reaction to overt negativity, "Sensitivity to potential harm"). Meanwhile, when going to four clusters, more nuance enters in (e.g., "Context and intent matter"). In other words, it seems that 1) overall sensitivity to toxicity is an important dimension in explaining variation, and 2) there are clusters of people who hold more nuanced views. For OpinionQA, descriptors that have to do with politics are often used (e.g., "Economically Conservative, but Populist on Trade"). For HK, which required more clusters to effectively predict, there are many specific phrases about what kinds of things the rater may or may not find offensive (e.g., "Profanity tolerance"; "Discomfort with stereotyping"; etc.). Meanwhile, for PR the clusters center around potentially conflicting chatbot preferences (e.g., "Appreciates simplicity" vs. "Appreciates nuanced and comprehensive answers"). Because value profiles are interpretable (see §7) and can recover demographic groupings where predictive, we have reason to believe that these qualitative differences map to important dimensions of disagreement for a dataset.
7 Extrinsic Evaluation
Value profiles are interpretable. We first explore interpretability – i.e., do the value descriptions change the decoder outputs in a common-sense manner? Because the encoder is prompted and only the decoder is trained (cf. Section 2.2), we believe that this serves as a strong regularization so that the value profiles correlate with held-out ratings only by the natural language values described. To ensure that this is the case, we test the interpretability of the autoencoder as follows: 1. For each instance and 100 value profiles, we get the estimated output distribution for the decoder. 2. We select the value profiles that have the largest Jensen-Shannon divergence. 3. We create a binary classification task: given an instance and two value profiles, which profile corresponds to which estimated output distribution? If the change in distribution is not correlated with common sense, we would expect 50% accuracy, while performance would approach 100% if an observer is always able to match value profiles to corresponding distributions.
Data Accuracy (%)
OQA 94.8 ($\pm 2.5$)
HK 96.3 ($\pm.6$)
DIC 95.5 ($\pm.7$)
VP 91.7 ($\pm.9$)
HL 89.8 ($\pm.9$)
PR 80.0 ($\pm.3$)
Chance 50
Table 2: Profiles are semantically interpretable (95% CI).
As an example of the task, here is a shortened example from VP: Is it moral, immoral, or morally dependent on context to do this: “Choosing not to get a vaccine." Profile 1: Prioritization of collective good over individual needs. Profile 2: Strong belief in individual liberty and autonomy. X probabilities: Moral: 92%, Context-dependent: 7%, Immoral: 2%. Y probabilities: Immoral: 94%, Context-dependent: 3%, Moral: 3%. Which profile goes with the X probabilities? Correct answer in footnote.  [^5] Profile two is the correct answer.
We report accuracies for a zero-shot prompted Gemini in Table 2. Accuracies range from 80-96% across all datasets, demonstrating that variation in outputs from value profiles are explainable by their plain natural language (p<<.001).
Figure 9: Calibration plots for value profile decoders. (Perfect calibration == dotted line). The decoders are very well-calibrated.
Decoders are well-calibrated. Decoder calibration is important for two principal reasons. Firstly, an appropriately calibrated decoder allows us to trust the model confidence w.r.t. error rate. Secondly, even raters with shared values may have varied outputs - a well-trained decoder would model this distribution appropriately. Calibration plots for the value profile decoders can be found in Figure 9. The trained decoders are quite well-calibrated, suggesting that we can generally trust the decoder’s output confidence.
Figure 10: Instance-level observed vs. estimated inter-annotator agreement (as the probability that two raters agree). The predicted simulated agreement correlates with the observed agreement.
Simulating an annotator population with value profiles. Given a trained decoder and a set of value profiles, one can simulate a Given a trained decoder and a set of value profiles, one can simulate a population – or “jury" (Gordon et al., 2022) – of annotators on novel instances. While one can do many things with such a simulated population (Park et al., 2023; Gordon et al., 2022), one experiment is to predict which instances raters would have higher or lower inter-annotator agreement (IAA).
In order to calculate out-of-distribution IAA, we first eliminate the datasets where annotators have no overlap (PR) and for which all raters annotated the same instances (DIC, OQA). For each instance, we then sample 100 value profiles that were not fit on that instance, and calculate the estimated probability of agreement between those annotators (assuming each rater annotates at random from the decoder’s output distribution). We also filter to instances that were labeled by a minimum number of annotators (See §A for details). We then compare to the actual observed probability that two raters agree on the instance (see Figure 10). For all three datasets, there is a positive correlation between the estimated and observed IAA, and this correlation is significant ($p<.001$) for HK and VP. While not much variance is explained ($R^{2}<.2$), the observed P(agree) is a high variance estimate with few ($\sim 5$) raters per instance. In summary, a simulated population with value profiles provides some explanatory power at predicting inter-annotator agreement, but is not yet a high precision tool.
8 Related Work
Clustering and demographics While aligning to groups can increase agreement (Chen et al., 2024), it also has been shown to flatten intra-group variation (Orlikowski et al., 2023; Wang et al., 2025) lead to stereotyping (Cheng et al., 2023), or simply not be correlated with subjective NLP tasks Orlikowski et al. (2025). Prior work explores embedding-based methods for clustering individuals by responses (Vitsakis et al., 2024; Li et al., 2024) and similarly finds that clusters cut across demographic groups. Beyond predictivity, demographics can still be important to collect for evaluating group fairness (Aguirre et al., 2023; Kirk et al., 2024b).
Steering to individuals Prompted large language models (LLMs) have been used to simulate human judgments, e.g.: NLP task annotators (Bavaresco et al., 2024), political survey respondents (Argyle et al., 2023), fact-checking labels (De et al., 2024), or human attitudes and behavior (Park et al., 2024). Textual user profiles have also been proposed for personalizing chats (Zhang et al., 2018). Hu and Collier (2024) similarly use textual demographic descriptions and find that they provide small, but statistically significant, gains in explaining human variation. Many works focus merely on prompting an existing LLM, while our work explicitly trains an LLM to better match varied perspectives (as in Gordon et al. 2022; Jiang et al. 2024). Encoding individual information from demonstrations is also analogous to behavioral user modeling for recommender systems (Radlinski et al., 2022; Ramos et al., 2024). Poddar et al. (2024) also use an autoencoder to model human variation, but focus on preference data and use a vector-valued latent space instead of natural language.
Values and alignment Similarly to natural language value profiles, Bai et al. (2022)’s "constitutional AI" train models to follow textual principles, although they focus on a single set of principles. Findeis et al. (2024) propose to learn preference principles directly from preference data (similar to our encoder setup). Values have also been normatively proposed as an alignment target (Gabriel, 2020; Klingefjord et al., 2024), and pluralistic alignment (Sorensen et al., 2024b) seeks to align AI systems to diverse values.
9 Conclusion
@@ Acknowledgements
10 Limitations
We have tried to test for generalization across six tasks and datasets and more than twenty demographic distributions, However, all of the experiments use the Gemma-2 (Gemma Team et al., 2024) and Gemini (Team et al., 2024) families of models. This is due in part due to the TPU hardware available to us and because of the expense of the experiments (more than 650 training runs). We have no reason to think that our results are due to anything particular about these families of models though, and prior work doing similar experiments on demographics with other models has reached similar results (Orlikowski et al., 2023; Hwang et al., 2023). That being said, future work could benefit from experiments on more model families.
11 Ethical Considerations
Profiling risks One of the potential ethical risks of the value profile through an autoencoder paradigm is in the name: “profile." Value profiles are, inherently, guesses about the underlying values that people may hold that lead them to annotate in the way that they do. There are potential privacy concerns here - people may not wish to have their underlying values exposed (Tomasev et al., 2021). It may be better if people had agency to create their own value profiles through some voluntary elicitation process (Park et al., 2024).
False generalization On the one hand, value profiles are an attempt to reduce the (often false) generalization inherent when grouping e.g. by demographic groups Dev et al. (2022), improving on widely used current techniques. On the other hand, generalization risks remain. For example, there may be multiple possible underlying values that could support a set of rater annotations. In the absence of additional information, the value encoder may arbitrarily assign a guess to the underlying values. Also, our experiments focus on English-language value profiles, so generalization to other languages is unknown. Additionally, there is always a risk of misrepresentation when using simulated human ratings in place of actual ratings at all (Agnew et al., 2024).
Interpretability, understandability, and user agency However, there are also several positive attributes to value profiles. First of all, they are interpretable - and therefore, potentially more understandable to a user. While people interact with many technologies today that are trying to model their behavior and preferences, most such systems do not break down their user representation into a format that as as easy to understand as a textual description. Additionally, this makes value profiles intervenable - people could change how they choose to be represented Balog et al. (2019); Lazar et al. (2024). Relatedly, value profiles serve as a step towards explainable AI (Arrieta et al., 2019; Koh et al., 2020) for human variation.
Enabling value reflection Learning values from data, while allowing users to modify the values, is loosely related to John Rawls’ concept of reflective equilibrium (Rawls, 2005; Knight, 2025): ratings are akin to judgments, and value profiles are an attempt to draw general “principles" out of the judgments in a bottom-up manner. Meanwhile, a user can then edit the value profile, applying top-down reflection on whether the values/principles are ones that the person would endorse. In this light, perhaps “value profiles" could help a person to explore their own value system, both in the values that their decisions may imply and considering which values they would reflexively endorse.
"Chosenness" of values Many works modeling diversity focus on socio-demographics. However, many demographics are not a result of a person’s agency, but rather a product of unchosen life factors - for example, the country in which one is born, or the economic opportunities available to them. Meanwhile, while the values that one holds can certainly be affected by unchosen factors (Nguyen, 2024), values can also be chosen for oneself. Thus, in the spirit of luck egalitarianism (Dworkin, 2002) it may be more justifiable to represent someone using values that they reflexively endorse, as opposed to boxing them in to the characteristics of a group they may not have chosen.